OpenStudy (anonymous):

Add and simplify if possible:

OpenStudy (anonymous):

\[3m \over 5 \] + \[2m \over 5\]

OpenStudy (chaise):

\[\frac{5m}{5}\]

OpenStudy (anonymous):

Answer choices are: A. 5m/10 B. m/2 C. m

OpenStudy (anonymous):

\[\frac{3m}{5}+\frac{2m}{5}=\frac{3m+2m}{5}=\frac{5m}{5}=m\]